Perform the indicated operation.

-16-(-14)
_________
-18-(-18)

Answer by edjones(8007)

-16-(-14)=-16+14=-2
-18-(-18)=-18+18=0
.
-2/0= answer is undefined. division by zero is not allowed.
.
Ed